OpenStudy (anonymous):

Find the following product: -3x^2y(-2x^5y^7) A. 6x^10y^7 B. -5x^7y^8 C. 6x^7y^8 D. -3x^2y-2x^5y^7

jimthompson5910 (jim_thompson5910):

Multiply the coefficients -3 and -2 to get (-3)*(-2) = 6. So the coefficient in the final answer is 6 Now add the corresponding exponents. So for x, add 2 an 5 to get 2+5 = 7. So the exponent for x in the final answer is 7 Do the same for y to get 1+7 = 8. So the exponent for y in the final answer is 8 So this means that the answer is \[\large 6x^7y^8\]

OpenStudy (anonymous):

c

OpenStudy (anonymous):

??

jimthompson5910 (jim_thompson5910):

yep